ID No Page 8 oi 8 this oints) Get the maximum likelihood estimator of θ and compute the estimate f data.
Answer #1

Here we have

f(x_{1})= heta (x_{1})^{ heta-1}

f(x_{2})= heta (x_{2})^{ heta-1}

.

.

.

f(x_{n})= heta (x_{n})^{ heta-1}

So the likelihood function is

0-1 71

Taking log of both sides gives

lnleft [L( heta) ight ]=nlnleft [ heta ight ]+( heta-1) sum_{i=}^{n}ln(x_{i})

Differentiating both sides gives

i=1

Equating it equal to zero gives

2 n(i) 0

7L :=1 In(.r.)

Hence, required estimate is

7L :=1 In(.r.)